The Q Company is carrying $1,000,000 worth of debt at a cost of debt (rd) of 5%. Q has 100% payout ratio and has zero growth expected. It’s current EBIT is $1 million, and its tax rate is 40%, and its cost of equity (rs) is 7%. The firm has 1,000,000 shares of common stock.

What is Q’s current stock price?

The CFO is currently evaluating a project which would have the firm buy back the old debt by issuing $1.5 million worth of new debt to replace the old debt and buy back stock. The new debt would have a cost of 8% and would increase the cost of equity to 10%.

What is the new stock price? And should they do the buy back?
Answer #1

Answer:-

Debt amount = $ 1000000
Coat of debt = 5 %
payout ratio =100 %
Growth rate = 0 %

Given EBIT = $ 1 Million = $ 1000000
Tax rate = 40 %
Coat of equity = 7 %
Number of shares outstanding = 1000000
Current stock price = ?
EBIT = $ 1000000
Interest expense = Debt x cost of debt = $ 1000000 x 0.05 = $ 50000
EBIT - Interest Expense = EBT
EBT = $ 1000000 - $ 50000
EBT = $ 950000
Net Income or PAT = EBT - Taxes
Taxes = $ 950000 x 0.40 = $ 380000
Net income or PAT or Earnings = $ 950000 - $ 380000 = $ 570000
Dividends / share = $ 570000 / 1000000 ( since there is 100 % payout ratio)
Dividends / share = 0.57

Dividends = $ 0.57 / share
Value of stock = D x ( 1+g) / (r-g)
Value of stock = $ 0.57 x 1 / (0.07 - 0) [ Given r = 7 % = 0.07 & growth g = 0 )
Value of stock = $ 8.14
Q's current price of stock = $ 8.14

The project which would have the firm buy back the old debt by issuing $1.5 million worth of new debt to replace the old debt and buy back stock.

Cost of new debt = 8 %
New debt will buy back the old debt  and buy stock
Interest expense = debt x cost of debt = $ 1500000 x 0.08 = $ 120000
EBIT = $ 1000000
EBT = $ 1000000 - $ 120000 = $ 880000
Taxes = $ 880000 x 0.40 = $ 352000
Net income or Earnings = EBT - Taxes = $ 880000 - $ 352000 = $ 528000
The number of shares left after buyout of shares worth $ 0.5 million shares
Number of shares bought back = $ 0.5 million / $ 8.14
= $ 500000 / $ 8.14
= 61425 ( approximately)
Number of shares outstanding after buyback = 1000000 - 61425 = 938575

New Stock price of Q = D x ( 1+g) / (r-g)
Considering 100 % payout ratio [ earnings = dividends]
Dividend / share = $ 528000 / 938575 = $ 0.5625
New Stock price of Q = $ 0.5625  ( 1+g) / 0.10 - g) [ r =10 % & g = 0 %]
New Stock price of Q = 0.5625 / 0.10
New Stock price of Q = $ 5.625

The buy back is not advisable as the cost of equity is high and therefore the stock price decreases from $ 8.14 to $ 5.625.